OpenStudy (anonymous):

how do you analyze this ? f(x)=x^2+1/x

OpenStudy (abb0t):

Graph it.

OpenStudy (anonymous):

yep ^

OpenStudy (anonymous):

i did @abb0t

OpenStudy (anonymous):

try using this? http://www.wolframalpha.com/input/?i=analyze+f%28x%29%3Dx%5E2%2B1%2Fx

OpenStudy (anonymous):

what did you notice about the function? @Ephilo

OpenStudy (anonymous):

but the question says to analyze and graph

OpenStudy (anonymous):

\[\frac{x^2+1}{x}= \frac{x^2}{x}+\frac{1}{x}= x+\frac{1}{x}\] no idea waht you mean by analyze

OpenStudy (anonymous):

well in analyzing you can include that you noticed and analyzed based in your graph the range and domain. Also what type of graph it is. Its min and max. Stuff like that. :)

OpenStudy (anonymous):

thanks everyone !:D